Ziggy's Famous Yogurt blends regular yogurt that is 3% fat with its no-fat yogurt to obtain low-fat yogurt that is 1% fat. How many pounds of regular yogurt and how many pounds of no-fat yogurt should be mixed to obtain 60 pounds of low-fat yogurt?
Answer by ptaylor(2198)

Let x=amount of 3% yogurt needed
Then 60-x=amount of no-fat yogurt needed
Now we know that the amount of fat in the 3% yogurt (0.03x) plus the amount of fat in the no-fat yogurt(0) has to equal the amount of fat in the final mixture
(0.01*60). So our equation to solve is:
0.03x=0.01*60 or
0.03x=0.6 divide both sides by 0.03
x=20 lb --------------amount of 3% yogurt needed
60-x=60-20=40 lb ----amount of no-fat yogurt needed
Ck
0.03*20=0.01*60
0.6=0.6
Hope this helps---ptaylor